OpenStudy (anonymous):

for t>=0, find when tlnt-t>0

OpenStudy (anonymous):

u want to \[t \ln t-t >0\]factor out \(t\) u have\[t(\ln t-1)>0\]now note that \(t>0\) so u just need to find \(t\) such that \(\ln t-1>0\)

OpenStudy (kropot72):

Following on from the explanation by @mukushla: \[\ln t >1\] Consider what happens when ln t = 1. In this case t equals the exponential number e. Therefore the required condition (t * ln t) - t >0 is met when t > e

OpenStudy (kropot72):

@omgitsana I hope you now understand the solution.